OpenStudy (anonymous):

Σ ((cos x)^2)/(n^2 + 1) from 1 to infinity Σ ((cos x)^2)/(n^2 + 1) from 1 to infinity @Mathematics

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \cos ^{2} x /n ^{2} + 1\]

OpenStudy (anonymous):

Are you asking what that sum is, or whether or not it converges?

OpenStudy (anonymous):

From Mathematica 8.04:\[\sum _{n=1}^{\text{Infinity}} \frac{\cos ^2(x)}{n^2+1}\text{=}\frac{1}{2} (\pi \coth (\pi )-1) \cos ^2(x)\]

OpenStudy (anonymous):

The actual sum is pretty gross, so I'll assume you mean to ask if it converges. The comparison test says that if we have two series a and b where for all n, \[a_n \geq b_n \] and a converges, then b must also converge. We can use this by noting that, since \[ cos^2(x) \leq 1 \] \[\frac{cos^2(x)}{n^2+1} \leq \frac{1}{n^2+1} \] and furthermore, since \[ n^2+1 > n^2 \] we have that \[ \frac{cos^2(x)}{n^2+1} \leq \frac{1}{n^2+1} < \frac{1}{n^2} \] and since the series \[\sum\frac{1}{n^2}\] converges, then our series converges too.

OpenStudy (anonymous):

awesome thank you! but how do you know that 1/n^2 converges?

OpenStudy (anonymous):

That is a pretty common series called a p-series. It turns out that \[\sum\frac{1}{n^p} \] converges for all p greater than or equal to 1. You can prove this via the integral test, which says that if you have some sequence a, and you convert that sequence to a function, and that function's integral converges, then the sequence converges too. For example, \[ a_n = \frac{1}{n^p} \] let \[ f(x) = \frac{1}{x^p} \] assuming \[p \geq 1\] we have that \[ \int_1^{\infty} \frac{1}{x^p} dx = -(p-1)\frac{1}{x^{p-1}}|^{x=\infty}_{x=1} = p-1\] thus the integral converges, and the corresponding sequence also converges, if \[p \geq 1\]